OpenStudy (anonymous):

There are two consecutive odd numbers such that twice the smaller number plus one is ten more than the greater number. What is the smaller number? anyone understand this!?

OpenStudy (anonymous):

2a+1=b-10

ganeshie8 (ganeshie8):

Let n, n+2 be two consecutive odd numbers now, we can express the problem as , 2n + 1 = (n+2) + 10

OpenStudy (anonymous):

the answer is 1

OpenStudy (anonymous):

My answers are 7, 9, 11, & 13

ganeshie8 (ganeshie8):

11